OpenStudy (anonymous):

Simplify. 2 to the 3 power + 3 to the 4 power

OpenStudy (anonymous):

2*2*2+3*3*3*3=?

OpenStudy (anonymous):

i got 89

OpenStudy (anonymous):

2*2*2+3*3*3*3=8+81=89